Raindrops fall vertically at 4m/s relative to the Earth. A train passenger observes the raindrops make an angle of 20 degrees to the vertical when the train is moving horizontally at constant speed. Calculate how fast the train is moving. Express your final answer in meters per second.


1
Expert's answer
2021-07-25T09:23:04-0400

"V_r=4km\/sec"

"\theta=20°"

"tan\theta=\frac{v_t}{v_r}"

"v_t=v_r tan\theta"

"v_t=4\times tan20°=4\times
\\v_r=4\times0.3639=1.45m/esc"